### v2

@@ I INTRODUCTION
Second, we present a generalized coordinate representation of the Cartesian stiffness forces from the DER theory, which is better suited for integration into MuJoCo and reduces computational load. In addition, the adoption of DER theory’s quasistatic treatment of the rod’s material frame twist improves stability and minimizes unnatural oscillations of the DLO simulation. Following that, we conduct two validation tests presented in the original literature [^1] Video of simulation validation for adapted model: https://youtu.be/-C48201PiUA [9] on our model and MuJoCo’s native cable model. Details can be found in Section III-D and Section IV.
@@ IV-A Simulation speeds
The increase in computational time for adapted is small ($2-4\%$) as compared to direct ($8-15\%$). This difference is attributed to adapted bypassing the $n+1$ distinct queries for the Jacobian at each simulation step. It is worth noting that the computational time increase for native ($1-3\%$) is smaller than adapted as the latter goes through a more complex stiffness computation.
@@ IV-B Validation
To guarantee accuracy of adapted and native, this section evaluates each model against available analytical solutions for elastic rods in the localized helical buckling [27] and Michell’s instability tests [28]. direct is excluded from further evaluations as it has been tested to exhibit identical behavior to adapted. Videos of the simulated validation (including unstable simulation of the native cable model) are available at: https://youtu.be/-C48201PiUA.
@@
By introducing and increasing twist along the tangent axis to an elastic rod loop, buckling occurs at a critical twist angle which can be determined analytically as $\theta^{n}_{c}=2\pi\sqrt{3}/(\beta/\alpha)$ when the loop radius $R=1$ [29]. This phenomenon shows the coupling between rod twist and equilibrium configuration. Fig. 4 shows the graph of critical buckling angle against $\beta/\alpha$ that was obtained for each simulation model. The average error defined as the 2-norm error per data point for adapted of 0.0643 is almost an order of magnitude smaller than native with 0.5950. adapted provides results significantly closer to the analytical solution. Data of native seems to exhibit a linear relation between $\beta/\alpha$ and $\theta^{n}_{c}$, which is inconsistent with a theoretical elastic rod. In Section VII, we will discuss in detail the possible shortcomings of native that causes its failure and how adapted accounts for these limitations.
@@ VI REAL EXPERIMENTS
Figure 6: Normalized position error between the real and simulated environments for 4 different poses. The adapted model generally performs better than the native one. Spikes in position error for the black DLO in pose 2 and the red in pose 4 are caused by unnatural twist wave oscillations experienced by the native model. Snapshots of the experiments can be found in the video.
From Fig. 6, we observe that the adapted model is generally better at shape prediction for all DLOs. The normalized position error, calculated as the average 2-norm euclidean error of 11 ordered points, between real and simulated results was the smallest for the white DLO across all poses. This is likely due to it exhibiting greater elasticity than the others. Interestingly, we observed that cases of spiked position error (black in pose 2 and red in pose 4) only occurred for the native model. These large inaccuracies arose because of the unnatural twist wave oscillations in the native simulation model, where twist waves repeatedly travel back and forth along the DLO length. Without additional training from real data, our parameter identification pipeline along with accurate simulation was able to consistently predict shapes with position errors per node of less than $5\%$ of total length.

### v3

@@ I INTRODUCTION
Second, we present a generalized coordinate representation of the Cartesian stiffness forces from the DER theory, which is better suited for integration into MuJoCo and reduces computational load. In addition, the adoption of DER theory’s quasistatic treatment of the rod’s material frame twist improves stability and minimizes unnatural oscillations of the DLO simulation. Following that, we conduct two validation tests presented in the original literature [^1] Video of simulation validation for adapted model: https://youtu.be/QY0B5IChghw [9] on our model and MuJoCo’s native cable model. Details can be found in Section III-D and Section IV.
@@ IV-A Simulation speeds
The increase in computational time for adapted is small ($-1\%\text{ to }3\%$) as compared to direct ($2\%\text{ to }15\%$). This difference is attributed to adapted bypassing the $n+1$ distinct queries for the Jacobian at each simulation step. It is worth noting that the computational time increase for native ($-3\%\text{ to }2\%$) is smaller than adapted as the latter goes through a more complex stiffness computation. A negative percent for some results could be attributed to the plain model going through a more computationally expensive time step due to its less structure movement.
@@ IV-B Validation
To guarantee accuracy of adapted and native, this section evaluates each model against available analytical solutions for elastic rods in the localized helical buckling [27] and Michell’s instability tests [28]. direct is excluded from further evaluations as it has been tested to exhibit identical behavior to adapted. Videos of the simulated validation (including unstable simulation of the native cable model) are available at: https://youtu.be/QY0B5IChghw.
@@
By introducing and increasing twist along the tangent axis to an elastic rod loop, buckling occurs at a critical twist angle which can be determined analytically as $\theta^{n}_{c}=2\pi\sqrt{3}/(\beta/\alpha)$ when the loop radius $R=1$ [29]. This phenomenon shows the coupling between rod twist and equilibrium configuration. Fig. 4 shows the graph of critical buckling angle against $\beta/\alpha$ that was obtained for each simulation model. The average error defined as the 2-norm error per data point for adapted of 0.0483 is almost an order of magnitude smaller than native with 0.7725. adapted provides results significantly closer to the analytical solution. Data of native seems to exhibit a linear relation between $\beta/\alpha$ and $\theta^{n}_{c}$, which is inconsistent with a theoretical elastic rod. In Section VII, we will discuss in detail the possible shortcomings of native that causes its failure and how adapted accounts for these limitations.
@@ VI REAL EXPERIMENTS
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionAB
real
adapted
native
Figure 6: a: Normalized position error between the real and simulated environments for 4 different poses. The adapted model generally performs better than the native one. b: Significant shape differences in the same real experiments for the black wire in pose 2 (top row – deformation of the native model protrudes above the horizontal plane from the grasp points) and the red wire in pose 4 (bottom row – loop formed in the native model occurs near the middle rather than the end of the wire). For these cases, the adapted model (center column) performed better than the native model (right column) in terms of shape comparison to real wires (left column).
From Fig. 6, we observe that the adapted model is generally better at shape prediction for all DLOs. The normalized position error, calculated as the average 2-norm euclidean error of 11 ordered points, between real and simulated results was the smallest for the white DLO across all poses. This is likely due to it exhibiting greater elasticity than the others. Interestingly, we observed that cases of spiked position error (black in pose 2 and red in pose 4 shown in Fig. 6(b)) only occurred for the native model. These large inaccuracies arose because of the unnatural twist wave oscillations in the native simulation model, where twist waves repeatedly travel back and forth along the DLO length. Without additional training from real data, our parameter identification pipeline along with accurate simulation was able to consistently predict shapes with position errors per node of $<5\%$ of total length.